# KeyNode-Driven Geometry Coding for Real-World Scanned Human Dynamic Mesh Compression

### v1

@@ IV-B KeyNode-based Compression Hyper-parameters
The number $N^{res}_{b}$ of quantization levels for residuals is set to 128 for all sequences, and $N^{RT}_{b}$ for key node rotations and translations is set to 64. NCOC mode is used only for Soldier, Thomas, and Mitch because of their sufficiently accurate P-frame prediction. We assigned 100 key nodes to Soldier, Thomas, and Levi, 200 to Mitch, and 300 to Longdress, Basketball, and Dancer. More key nodes are assigned to sequences with more complex motion to enhance P-frame prediction.
@@ IV-C Analysis of KeyNode-based Encoded Components
This section offers a closer examination of our codec elements. For a dynamic mesh, we transmit the statically encoded I-frames, encoded key nodes, encoded RT, and encoded residuals for P-frames. Figs. 8 and 9 illustrate the proportion of each encoded component as the GoF size and number of key nodes vary. The proportion of encoded residuals increases the most with GoF size increases; this is due to a smaller number of I-frames for large GoF cases. When the number of key nodes increases, the encoded RT component increases proportionally, while the geometry residual proportion decreases, because a larger number of key nodes does better P-frame prediction, leading to lower prediction errors.
@@
Figure 9: Encoded components when varying number of key nodes in Thomas. GoF size is 3 frames.
IV-D Distributions of Encoding parameters
In our codec, RT and residuals are quantized and Huffman-coded with a fitted Cauchy distribution. Fig. 10 presents the distributions of our data with the fitted Gaussian, Laplace, and Cauchy distributions. The Cauchy distribution has the best alignment with the data. Y-axis and Z-axis residuals have a similar histogram with the X-axis in Fig. 10a, and key node translations have a similar histogram with Fig. 10b.
Figure 10: Distributions of Residuals and RT in frame 618 of Thomas. The histograms are plotted with 200 bins.
IV-E Geometry Errors at each Compression Stage
@@
IV-F Influence of GoF Size and I-frame Quality
In this section, we evaluate various GoF sizes and I-frame qualities to understand how they impact compression performance. The experiment is done on Thomas’s first ten frames, with four different GoF sizes and four different I-frame quality levels. We evaluate the I-frame quality levels 1 to 4 from lower to higher. For P-frames, we use the same encoding configuration across all cases being assessed.
Fig. 12 shows that smaller GoF sizes are more sensitive to variations in I-frame quality, as indicated by a wider spread in compression performance when changing quality levels. Smaller GoF sizes lead to higher bit rates, whereas larger GoF sizes tend to perform better at lower bit rates. For instance, at similar quality settings, GoF size 10 achieves a lower bit rate. Overall, Fig. 12 suggests that the KeyNode-based method prefers medium to high I-frame quality. When aiming for lower target bit rates, it is advantageous to increase the GoF size rather than reduce the quality of the I-frames.
Figure 12: Influence of GoF size and I-frame quality on compression performance - first ten frames in Thomas sequence.
@@ IV-F Influence of GoF Size and I-frame Quality
IV-G Influence of P-frame Encoding Components
@@
Figure 13: Influence of P-frame compression parameters for the first ten frames of Basketball with GoF size = 5.
It is evident that the octree size for residual coding significantly impacts the rate, as the squares are predominantly on the right side, while the circles are clustered on the left. With a notable increase in rate, enlarging the octree usually leads to an improvement in quality, as the residuals are described more accurately with bigger octrees.
In Fig. 13, similar hatch patterns tend to form distinct vertical clusters, suggesting that $N^{res}_{b}$ strongly influences the rate.
Unlike the octree size and $N^{res}_{b}$, varying $N^{RT}_{b}$ does not strongly impact the rate. However, those parameters primarily control the quality when the other two parameters are fixed.
Additionally, each value of $N^{RT}_{b}$ can cover the entire range on the rate axis, but the green and blue markers generally achieve lower distortion values with only a small increase in rate. Consequently, it appears desirable to use larger values of $N^{RT}_{b}$.
IV-H Compression Performance Comparison
In this section, we compare the Rate-Distortion (RD) performance of our KeyNode-based method with the state-of-the-art V-DMC’s geometry coding method [21] with their two modes: all intra-coding (denoted by V-DMC - intra), and with inter coding frames (denoted by V-DMC - inter). V-DMC was chosen as the baseline because it demonstrated the highest effectiveness among methods responding to the MPEG call for proposals and offers a publicly available implementation, enabling fair and reproducible comparisons.
@@ IV-H Compression Performance Comparison
For the Fast Motion and Join/Split categories, Fig. 15 depicts the RD curves using MSDM2. Our methods are slightly better than V-DMC on the Longdress, Levi, and Basketball sequences and slightly worse on Dancer.
@@
The curves in Figures 14 and 15 can be summarized in terms of average bitrate savings using Bjøntegaard Delta rate (BD-rate) percentage, following the methodology in [35]. Table II presents the BD-rate percentages compared to two coding modes of V-DMC. BD-rate measures the difference in bit rate required to achieve the same level of distortion between two compression methods. A lower BD-rate indicates better compression efficiency, as it achieves the same quality with less data. Consequently, in Table II, negative values indicate that Ours outperforms the comparison method, while positive values indicate that Ours performs worse. According to Table II, our methods considerably outperform V-DMC for sequences in the Slow Motion Category. For sequences in the Fast Motion category, compared to V-DMC, our methods perform slightly better on Longdress, slightly worse on Dancer, and comparably on Levi. For the Join/Split category, we outperform V-DMC by a small margin.
@@
V-DMC inter FF -32.58 -46.12 -37.57 -23.05 9.61 0.12 -1.90 -18.78
ADP -37.14 -48.70 -48.99 -26.06 4.61 -13.45 -19.07 -26.97
V-DMC intra FF -36.71 -76.45 -41.78 -3.92 9.39 -5.70 -1.85 -22.43
ADP -40.45 -79.95 -50.91 -6.92 4.54 -17.56 -17.74 -29.86
Average -24.51
@@ IV-I Frame-wise and GoF-wise Distortion Analysis
Figure 16: Frame-wise (a) and GoF-wise (b) analysis for Thomas under similar sequence encoded size. The evaluated GoF size is 5 frames. Label format: Method (Sequence encoded size).
@@ IV-J Geometry Reconstruction Distortion
Figure 19: Reconstruction distortion of our method and V-DMC, under similar bitrate, for frames 618 to 622 in Thomas. The first mesh in Ours is an I-frame, and the four other meshes are forward P-frames.
Notably, with Thomas in Fig. 19, our method achieves lower errors with a similar bitrate. In the zoom-in region, our decoded mesh more closely resembles the original mesh than that of V-DMC, which tends to lose the details of the fingers.
@@
Figure 20: Frame 13 of levi, extracted from the decoded sequences using our method and V-DMC.
@@ V Discussion and Future Work
Our method presents notable advantages in geometry coding compared to V-DMC, particularly evident in the Slow Motion sequence category. For Fast Motion and Join/Split sequences, our approach demonstrates comparable performance. These results highlight the effectiveness of our method in handling varying topology and scan defects, while also emphasizing the challenge of accurately predicting the geometry of P-frames in human dynamic meshes with rapid motions.
Our KeyNode-driven codec introduces a new way to compress scanned 3D human dynamic models, opening up numerous avenues for future research and development. One promising direction is the adaptive placement of I-frames, where dynamically determining the optimal points for I-frame insertion could further enhance compression performance. Another potential improvement is the integration of view-dependent coding, which takes into account that users typically view a 3D model from a single angle at a time. By reducing the quality of unseen areas, we can lower bitrate requirements without compromising the user experience. Additionally, while the current KeyNode-driven codec delivers effective results, its encoding complexity remains a limitation. The optimization-based nature of components such as the Optimal Key Node Generator and Rotation and Translation Extractor can be computationally demanding. Future work could focus on relaxing these optimization steps and exploring parallelization techniques to improve efficiency. Furthermore, the KeyNode-driven codec holds great potential for broader applicability by adapting the prediction model to align with the characteristics of diverse mesh structures. While the current method is optimized for objects exhibiting a combination of rigid and non-rigid motions, such as dynamic human meshes, future work could explore alternative transformation prediction strategies or tailored strategies for different types, unlocking its versatility across a range of scanned models.

### v2

@@ IV-B KeyNode-based Compression Hyper-parameters
The number $N^{res}_{b}$ of quantization levels for residuals is varied across $2^{7},2^{8}$, and $2^{9}$. The codebook size $N^{RT}_{b}$ for key node rotations and translations is varied across $2^{3},2^{5}$, and $2^{7}$. The size of the unbalanced octree for Residual coding varies across 20, 100, and 500 leaf nodes. The GoF size varies across 2, 4, 8, 16, and 32. We assigned 100 key nodes to Soldier, Thomas, Mitch, and Levi, and 300 to Dancer, Longdress, Basketball. More key nodes are assigned to sequences with more complex motion to enhance P-frame prediction.
@@ IV-C Analysis of KeyNode-based Encoded Components
This section offers a closer examination of our codec elements. For a dynamic mesh, we transmit the statically encoded I-frames, encoded key nodes, encoded RT, and encoded residuals for P-frames. Figs. 9 and 10 illustrate the proportion of each encoded component as the GoF size and number of key nodes vary. The proportion of encoded residuals increases the most with GoF size increases; this is due to a smaller number of I-frames for large GoF cases. When the number of key nodes increases, the encoded RT component increases proportionally, while the geometry residual proportion slightly decreases, because a larger number of key nodes does better P-frame prediction, leading to lower prediction errors.
@@
Figure 10: Encoded components when varying number of key nodes in Thomas. GoF size is 8 frames.
IV-D Geometry Errors at each Compression Stage
@@
IV-E Influence of GoF Size and I-frame Quality
In this section, we evaluate various GoF sizes and I-frame qualities to understand how they impact compression performance. The experiment is done on Thomas’s first 32 frames, with five different GoF sizes and four different I-frame quality levels. We evaluate the I-frame quality levels 1 to 4 from lower to higher. For P-frames, we use the same encoding configuration across all cases being assessed, and the prediction mode is Fully Forward.
Fig. 12 shows that smaller GoF sizes are more sensitive to variations in I-frame quality, as indicated by a wider spread in compression performance when changing quality levels. Smaller GoF sizes lead to higher bit rates, whereas larger GoF sizes tend to perform better at lower bit rates. For instance, at similar quality settings, GoF size 16 achieves a lower bit rate. Overall, Fig. 12 suggests that the KeyNode-based method prefers medium to high I-frame quality. When aiming for lower target bit rates, it is advantageous to increase the GoF size rather than reduce the quality of the I-frames.
Figure 12: Influence of GoF size and I-frame quality on compression performance - first 32 frames in Thomas sequence.
@@ IV-E Influence of GoF Size and I-frame Quality
IV-F Influence of P-frame Encoding Components
@@
Figure 13: Influence of P-frame compression parameters for the first ten frames of Mitch with GoF size = 4.
The dominant factor influencing overall quality is the precision of motion transformations, determined by the number of quantization bins $N^{RT}_{b}$. When $N^{RT}_{b}$ is low (e.g., $2^{3}$, orange markers), motion prediction becomes inaccurate, leading to high distortion. In this case, increasing the number of key nodes, refining octree size, or improving residual quantization cannot compensate for the poor motion modeling, highlighting the importance of precise motion representation.
Once $N^{RT}_{b}$ is sufficiently large (e.g., $2^{5}$ or $2^{7}$, green and blue markers), increasing the number of key nodes slightly improves distortion by refining motion granularity. Finer octree subdivisions can enhance adaptation to local residual structures, as observed in cases such as the 50-keynode green markers and 100-keynode orange markers with $N^{res}_{b}=2^{6}$ (cross-hatch pattern), where increasing the octree size from 5 to 20 leaf nodes reduces distortion. However, excessive partitioning—particularly in minimal residual scenarios such as the Mitch sequence—can sometimes disrupt residual coherence and amplify minor artifacts and unsmooth details, slightly degrading perceptual quality even though vertex positions improve in a least-squares sense. Finally, the influence of residual quantization $N^{res}_{b}$ is subtle when octree sizes are small but becomes increasingly pronounced as finer spatial subdivisions amplify the bitrate sensitivity to residual precision.
The performance difference between $N^{RT}_{b}=2^{5}$ and $2^{7}$ is small, suggesting that once a sufficient precision is reached for motion transformations, additional refinement offers diminishing returns.
IV-G Compression Performance Comparison
In this section, we compare the Rate-Distortion (RD) performance of our KeyNode-based method (under ADP mode) with the state-of-the-art V-DMC’s geometry coding method [21] with their two modes: all intra-coding (denoted by V-DMC - intra), and with inter coding frames (denoted by V-DMC - inter). V-DMC was chosen as the baseline because it demonstrated the highest effectiveness among methods responding to the MPEG call for proposals and offers a publicly available implementation, enabling fair and reproducible comparisons.
@@ IV-G Compression Performance Comparison
For the Fast Motion and Join/Split categories, Fig. 15 depicts the RD curves using MSDM2. Our method is slightly better than V-DMC on Longdress, and notably outperforms V-DMC on Dancer, Levi, and Basketball.
@@
The curves in Figures 14 and 15 can be summarized in terms of average bitrate savings using Bjøntegaard Delta rate (BD-rate) percentage, following the methodology in [35]. BD-rate measures the difference in bit rate required to achieve the same level of distortion between two compression methods. A lower BD-rate indicates better compression efficiency, as it achieves the same quality with less data. Table II presents the BD-rate percentages compared to two coding modes of V-DMC; negative values indicate that Ours outperforms the comparison method, while positive values indicate that Ours performs worse. According to Table II, our methods considerably outperform V-DMC for sequences in all categories.
@@
V-DMC inter -93.6 -91.55 -87.55 -36.78 -24.55 -58.57 -34.37 -60.99
V-DMC intra -77.3 -94.66 -82.97 -18.74 -23.62 -59.14 -34.71 -55.87
Average -87.94 -36.9 -34.54 -58.43
IV-H Ablation studies
We conducted a leave-one-out ablation study for three representative sequences (Soldier for slow-motion, Levi for fast-motion, and Basketball for join-split category), where we begin with the full method and disable one component at a time. This allows us to assess the individual contribution of each module by observing the performance drop when it is removed. The ablation, shown in Figure 16, includes:
  • Decimation: Replace our key node generator with mesh decimation.
  • Uniform Quantization: Replace our Cauchy-based quantization and coding with uniform quantization and explicit Huffman tables for encoding of Transformations RT.
  • Fully Forward: Replace adaptive dual-direction prediction with a fixed forward prediction.
  • w/o RC: Disable residual coding.
  • Ours: Full method with all components enabled.
Figure 16: Ablation study for three representative sequences for three categories.
Figures 16a and 16b show that for the Levi and Soldier sequences, our Optimal Key Node Generator achieves significantly lower distortion compared to Mesh Decimation. However, for Basketball in Fig. 16c, the improvement is less pronounced, with the curves overlapping. This is likely due to the use of a search step size (20-200 nodes per step) in our generator for computational efficiency, which yields near-optimal rather than globally optimal key nodes. A finer search step would likely enhance performance further. Additionally, the large number of key nodes selected for Basketball (300 vs. 100 for Levi and Soldier) may diminish the visible difference between methods, suggesting that Mesh Decimation may suffice in scenarios with a large key node count, thus reducing computational complexity. Overall, the Decimation curves emphasize the effectiveness of our KeyNode-driven codec, particularly in capturing temporal motion in slow-motion sequences, while our Optimal Key Node Generator excels at selecting the right set of key nodes for this purpose.
For all categories, the Uniform Quantization curves show that the bitrate needed is significantly higher without fitting into a Cauchy distribution to convey the Huffman tables implicitly.
The w/o RC curves indicate that residual coding is particularly effective in the slow-motion category. This may seem counterintuitive when compared to traditional 2D video coding, where residual coding typically yields greater improvements in fast-motion sequences. In 2D video, pixel-to-pixel correspondence is explicit, but in scanned meshes, this correspondence is not available, and our residual estimation method is less effective for complex-motion sequences. As a result, even with perfect residual correction, the quality improvement is either small or negative, and therefore, Residual Coding is disabled for multiple frames. This suggests that improving residual estimation in complex-motion scenarios could be a promising area for future work.
@@ IV-I Frame-wise and GoF-wise Distortion Analysis
Figure 17: Frame-wise (a) and GoF-wise (b) analysis for Thomas under similar sequence encoded size. The evaluated GoF size is 4 frames. Label format: Method (Sequence encoded size).
@@ IV-J Geometry Reconstruction Distortion
Figure 20: Reconstruction distortion of our method and V-DMC, under similar bitrate for Thomas.
Notably, with Thomas in Fig. 20, our method achieves lower errors with a similar bitrate. In the zoom-in region, our decoded mesh more closely resembles the original mesh than that of V-DMC, which tends to lose the details of the fingers, despite appearing smoother due to the increased number of vertices from heavy subdivision.
@@
Figure 21: Frame 145 of levi, extracted from the decoded sequences using our method and V-DMC.
IV-K Theoretical Complexity Analysis
In this section, we evaluate the theoretical complexity of our method and V-DMC. Since the coding of I-frames is identical in both methods, we focus on comparing the complexity of the inter-mode, which is used for encoding predictive frames. It is important to note that the complexity estimates are based on theoretical concepts outlined in related papers, and the real-world complexity may vary depending on the actual implementation.
Tables III and IV summarize the estimated computational complexity of V-DMC and our method, respectively, where $V$ is the number of vertices in the mesh. The analysis shows that both methods exhibit comparable theoretical complexity when the number of subdivided vertices in V-DMC satisfies $S\leq V$. However, in scenarios where $S>V$, its computational cost can grow significantly. In contrast, our method maintains a more stable complexity profile by directly operating at the vertex level.
TABLE III: V-DMC’s theoretical complexity. $B$: number of vertices in the base mesh; $S$: number of vertices in the subdivided mesh.
Component Best-case Worst-case
Pre-processing $O(V\log V)$ $O(V^{2})$
Base Mesh Reuse Decision $O(S\log S)$
Motion Field Encoder $O(B)$
Displacement Coding $O(S\log S)$ $O(S^{2})$
Total if $V\geq S$ $O(V\log V)$ $O(V^{2})$
if $V<S$ $O(S\log S)$ $O(S^{2})$
TABLE IV: Our theoretical complexity. $N$: number of key nodes.
Component Best-case Worst-case
Key Node Generation $O(V\log V)$
Transformation Extractor $O(V\log V)$
Key Node Encoder $O(N\log V)$
Transformation Encoder $O(N)$
Residual Coding $O(V)$ $O(V^{2})$
Total $O(V\log V)$ $O(V^{2})$
@@ V Discussion and Future Work
Our method presents notable advantages in geometry coding compared to V-DMC, particularly evident in the Slow Motion sequence category. For Fast Motion and Join/Split sequences, the improvements are smaller but still considerable. These results highlight the effectiveness of our method in modeling the motion while handling varying topology and scan defects. It also emphasizes the challenge of accurately predicting the geometry of P-frames in human dynamic meshes with rapid motions.
Our KeyNode-driven codec introduces a new way to compress scanned 3D human dynamic models, opening up numerous avenues for future research and development. One promising direction is the adaptive placement of I-frames, where dynamically determining the optimal points for I-frame insertion could further enhance compression performance. Another potential improvement is the integration of view-dependent coding, which takes into account that users typically view a 3D model from a single angle at a time. By reducing the quality of unseen areas, we can lower bitrate requirements without compromising the user experience. Additionally, while the current KeyNode-driven codec delivers effective results, its encoding complexity remains a limitation. The optimization-based nature of components such as the Optimal Key Node Generator and Rotation and Translation Extractor can be computationally demanding. Future work could focus on relaxing these optimization steps and exploring parallelization techniques to improve efficiency. Furthermore, the KeyNode-driven codec holds great potential for broader applicability by adapting the prediction model to align with the characteristics of diverse mesh structures. While the current method is optimized for objects exhibiting a combination of rigid and non-rigid motions, such as dynamic human meshes, future work could explore extending the transformation types, e.g., adding Scaling and Shearing in addition to Rotation and Translation, unlocking its versatility across a range of scanned models.